Let h be an exponentially-distributed random variable with the distribution function p- exp(-x) for x > 0 and ph = nction Ph 0 for a s 0. Derive the distribution function of its square root, Solution: 2y exp(-y2
Answer #1

here let Y=√x

therfore cdf of Y =F(Y)=P(Y<y)=P(√X <y)=P(X<y2)=\int_{0}^{y^{2}}f(x) dx =\int_{0}^{y^{2}}exp(-x) dx =-exp(-x)|y^20

F(y)=1-exp(-y2)

therefore distribution function of y: f(y)=(d/dy)*F(y)=(d/dy)*(1-exp(-y2))

f(y)=2yexp(-y2)